OpenStudy (anonymous):

find the next 3 number in the sequence 1) 2,-2,4,-4,6 . . . 2) -15, -10,0,15,35 . . . these get me really confused

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

is it -6

OpenStudy (anonymous):

7, -7, 8 -8?!

OpenStudy (anonymous):

for the first one?!! am i right? :(

OpenStudy (anonymous):

1) -6, 8, -8 2) 60, 90, 125

OpenStudy (turingtest):

looks like the pattern for the first is 2,-2,2+2,-(2+2),2+2+2,-(2+2+2) so we continue this way ,2+2+2+2,-(2+2+2+2),2+2+2+2+2 =4(2),-4(2),5(2)=8,-8,10

OpenStudy (anonymous):

aww i was almost right for the first one lol

OpenStudy (turingtest):

oh sorry I didn't realize we had stopped at 6...

OpenStudy (turingtest):

Stars Mars is correct

OpenStudy (anonymous):

u are right ayojenn but the problem asked for the next 3 :)

OpenStudy (anonymous):

yes i just saw that right now lol :(

OpenStudy (anonymous):

i got carried away when i kinda got the hang of it lol